OpenStudy (lgbasallote):

Another brain exercise :DDD One day your minimalist friend is visiting, and she presents you with the numbers 2, 3, 4 and 5, and the symbols = and +, and the following tantalizing challenge: It is possible to use each number and symbol once and only once and end up with a true equation; further, you cannot use any number or symbol not given. How can you do this?

OpenStudy (lgbasallote):

I just copied this problem from a website :D this problem is not mine. (i'm avoiding plagiarism)

OpenStudy (anonymous):

YOu don't have better things to do? :P

OpenStudy (lgbasallote):

well...this site has so few questions so i thought of entertaining bored people with these :D

OpenStudy (experimentx):

i guess, it's impossible. you have four number's and two symbols, and one of the number constructed must be two digit number and smallest two digit number you can have is 23 and the greatest possible number from + operator is 5+4=9 so mathematically, equation is impossible

OpenStudy (anonymous):

\[3^2=4+5 \]

OpenStudy (lgbasallote):

oh but it is possible... :D you just need to think outside the box...and fool you cant use symbol not given...there is a hidden ^ symbol in exponents :p

OpenStudy (experimentx):

good reply, never considered that possibility, other than that, i guess it's not possible.

OpenStudy (anonymous):

Igb, do you have solution other than mine?

OpenStudy (experimentx):

@Igbasallote .. but he's not writing any ... has he?

OpenStudy (lgbasallote):

oh wait....it seems fool was right! congrats :DDD

OpenStudy (lgbasallote):

i thought you cant use exponents but it seems you can :D

OpenStudy (anonymous):

YAY! :D

OpenStudy (lgbasallote):

do you have cute problems other than the one you posted fool? id like to try problems too :DDD

OpenStudy (anonymous):

I sure do, but I don't do riddles these days :(

OpenStudy (lgbasallote):

nahhh..just interesting algebra thingies are fun :)

OpenStudy (anonymous):

it is a lot easier: 3+4=5+2 :D

OpenStudy (lgbasallote):

no..there's inly one + actually.. 3^2 = 4 + 5 is the only logical answer :D

OpenStudy (anonymous):

Incorrect, '+' is used twice!

OpenStudy (anonymous):

oh.. my mistake :'(